Two coherent sources of different intensities send waves that interfere. The ratio of maximum to minimum intensity is 25. The intensity ratio of the sources is

 
 
 

Grade:12

1 Answers

Govind Sharma
3 years ago
Let a1a1 and a2a2 be amplitudes of the two waves. For maximum intensity
Imax=(a2+a2)2Imax=(a2+a2)^2
For minimum intensity
Imin=(a2−a2)2Imin=(a2−a2)^2
Given, ImaxImin=251=((a1+a2)2)(a1−a2)2Imax/Imin=25/1=((a1+a2)^2)/(a1−a2)^2
⇒a1+a2a1−a2=51⇒a1a2=32⇒a1+a2/a1−a2=5/1
                                                           ⇒a1/a2=3/2
I1/I2 = a1^2/a2^2 = 9/4
∴I1I2=a21a22=(32)4=94